OpenStudy (anonymous):

what is the square root of 21

OpenStudy (phi):

4*4=16 5*5= 25 so between 4 and 5

OpenStudy (barrelracer011):

4.582575695

OpenStudy (mathstudent55):

\(\sqrt{21} \) It can't be simplified, but you can find an approximation.

OpenStudy (mrnood):

try this on calculator type 2 type 1 type sqrt key type =

OpenStudy (anonymous):

use calculator :3